OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. y = 3x + 2 y = x + 12 A. (5, 17) B. (0, 12) C. (0, 2) D. (3, 2)

OpenStudy (anonymous):

@ParthKohli Already showed you how to do it. Go to your last problems and see the work.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 3x + 2 = x + 12 }\) \(\Large \color{purple}{\rightarrow 3x = x + 10 }\) \(\Large \color{purple}{\rightarrow 2x = 10 }\) \(\Large \color{purple}{\rightarrow x = 5 }\) y = x + 12. Clearly, \(\Large \color{purple}{\rightarrow y = 12 + 5 }\) \(\Large \color{purple}{\rightarrow y = 17 }\) @Romero thanks :), but he wants to check them.

Parth (parthkohli):

A. (5,17) is the answer.

OpenStudy (anonymous):

@ romero iknow he did.. but im making sure I HAVE THE CORRECT ANSWERS. !!!!!